A charge of -2.7 µC is fixed at the center of a compass. Two additional charges are fixed on the circle of the compass, which has a radius of 0.094 m. The charges on the circle are -4.4 µC at the position due north and +4.50 µC at the position due east. What are the magnitude and direction of the net electrostatic force acting on the charge at the center? Specify the direction relative to due east. magnitude N direction °
Answer #1

X-component of the net force

Fx=K|Q||Q1|/r2 =(9*109)(2.7*10-6)(4.5*10-6)/0.0942=12.3755 N

Y-Component of the net force

Fy=-K|Q||Q2|/r2 =-(9*109)(2.7*10-6)(4.4*10-6)/0.0942=-12.1005 N

Magnitude of net force

F=sqrt[12.37552+(-12.1005)2]

F=17.31 N

Direction

o=tan-1(-12.1005/12.3755) =-44.36o

o=44.36o south of east